Circular No. 5639

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1992 No. 2 G. S. Rossano, R. J. Rudy, R. C. Puetter, and D. K. Lynch, Space and

Environment Technology Center, The Aerospace Corporation, communicate: “Near-infrared observations of N Sgr 1992 No. 2 have been obtained using a cooled scanning germanium spectrometer with a 10′′ × 8′′ oval aperture on the 3-m Shane telescope of Lick Observatory. Spectroscopic observations (range 0.75-1.35 μm, resolution 300-550) show that the nearinfrared spectrum has changed considerably since August. The source excitation is now higher, as evidenced by a significant increase in the strength of He II lines. He I at 1083.0 nm has increased considerably and is now the dominant line of the spectrum. Neutral oxygen 844.6- and 1128.7-nm features excited by the Lyβ fluorescence mechanism are still present. Weak O I at 777.4 and 1316.4 nm is also detected. Previously observed lines of neutral nitrogen and carbon are now weak or absent. Several lines remain unidentified. The nova was observed on three nights over a 4-day interval on Oct. 7.2, 8.2, and 9.2 UT. A 5-percent decrease in brightness was observed over this time interval. A monochromatic magnitude of 9.4 at 1.25 μm was observed on Oct. 9.2.”

COMET HELIN-LAWRENCE (1992q)

Orbital elements from MPC 20909:

Epoch = 1993 Apr. 3.0 TT

T = 1993 Mar. 15.0512 TT  ω = 268.^{o}7932
e = 0.965523  Ω = 194.6696 }2000.0
q = 2.040011 AU  i = 106.8513
1992/93 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 25 0^{h}33.^{m}13 −44^{o}26.^{′}0 1.965 2.611 120.^{o}2 19.^{o}2 14.1
Nov. 4 0 09.22 −47 47.7 2.046 2.543 108.6 21.7 14.1
14 23 47.64 −50 02.2 2.147 2.478  97.5 23.3 14.1
24 23 30.02 −51 24.6 2.255 2.415  87.1 24.1 14.1
Dec. 4 23 17.13 −52 11.6 2.363 2.357  77.6 24.1 14.1
14 23 08.95 −52 37.9 2.461 2.302  69.1 23.6 14.1
24 23 05.06 −52 54.1 2.544 2.251  61.7 22.6 14.1
Jan. 3 23 04.97 −53 08.0 2.607 2.204  55.5 21.6 14.0
13 23 08.16 −53 25.1 2.647 2.163  50.7 20.6 14.0
23 23 14.31 −53 49.6 2.663 2.127  47.6 20.0 13.9

1992 October 19 Daniel W. E. Green